OpenStudy (anonymous):

What is the resulting ordered pair if the value of the independent variable is 2? f(x) = –5x + 4 A. (–6, 2) B. (2, 14) C. (2, –6) D. (6, 2)

OpenStudy (anonymous):

independent Variable = x therefore substitute x=2 in the equation

OpenStudy (anonymous):

***REMEMBER** (x,y)

OpenStudy (anonymous):

do i multiply the 2?

OpenStudy (anonymous):

f(2) = -(5x2) +4

OpenStudy (anonymous):

2,14 ? is it because of the 5*2=10+4=14 am i wrong?